Diffusion Morning Report 10/26/03

ted berg tberg at snf.stanford.edu
Mon Oct 27 07:03:42 PST 2003


Good Morning All,
    Everything looking pretty good this morning.  Still investigating  
Nitride issues .  Epi doesn't seem to have had any glitches since
 Floppy controller was removed Boards should be here this week.  Tystar 
and LTO are getting close to change.

Diffusion 	Date: 	10/27/2003 	 
Tube Name/Number 	Status 	Comments 	THICKNESS-Change at
Tylan 1 	Up 	  	 
Tylan 2 	Up 	  	 
Tylan 3 	Up 	  	 
Tylan 4 	Up 	  	 
Tylan 5-BBr3 	Up 	  	 
Tylan 6-POCl3 	Up 	  	 
Tylan 7-Dirty Ox 	Up 	  	 
Tylan 8-SiGe 	Up 	  	12.5/30
Tylan 9-Poly 	Up 	  	5.13/30
Tylan 10-Nitride 	yellow 	bubbles- Investigating 	6.5/30+
Tylan 11-BPSG-LTO 	Up 	  	24.9/30
Tylan 12-FGA 	Up 	  	 
Tystar 	Up 	  	28.7/30
Doped Poly 	GONE 	  	 
TEOS 	yellow 	  	 
Tylan 16 	N/A 	  	 
Burn Box 	Up 	  	 
ZYGO 	Up 	  	 
  	  	  	 
GAAS 22 	UP 	  	 
GAAS 23 	UP 	  	 
  	  	  	 
SiH4 for furnaces 	  	Auto changeover working well 	Currently 500- B(c4) 
on line


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20031027/60d6b544/attachment.html>


More information about the diffusion mailing list